OpenStudy (anonymous):

Find the equation of the plane which passes through O and is parallel to x-5y+z=7

OpenStudy (anonymous):

x=−z+5y+7

OpenStudy (anonymous):

hopefully this helped you out

OpenStudy (amistre64):

parallel things differ by a constant ..

OpenStudy (amistre64):

"solving" for x does not produce a different parallel equation of a plane :/

OpenStudy (anonymous):

so how do you solve it?

OpenStudy (amistre64):

at best: take the point the give you .. the origin; and apply the normal parts from the plane that they gave you to it

OpenStudy (amistre64):

or, ignore the constant in the stated plane, and plug in the point to solve for the constant of the parallel plane

OpenStudy (amistre64):

x-5y+z=7 x-5y+z=k 0-5(0)+0=k 0 = k therefore: x-5y+z=0

OpenStudy (anonymous):

so u substitue the 0 to the variables to solve for the plane?

OpenStudy (anonymous):

more like to solve for the constant number, and that would give out the parallel plane?

OpenStudy (amistre64):

parallel planes have the same basic parts .... just like parallel lines have the same basic parts ... they differ by a constant; that constant is what moves them up to down

OpenStudy (anonymous):

alright, Thanks

OpenStudy (amistre64):

given a normal (a,b,c) and a point (xo,yo,zo) we define a plane as: a(x-xo)+b(y-yo)+c(z-zo) = 0 ax-axo+by-byo+cz-czo = 0 ax +by + cz - (axo+byo+czo) = 0 ^^^^^^^^^^ the constant parallel planes have the same abc parts, and the same generaic xyz parts; they differ only by the point used - which only alters the constant